shift summary, owl 02/05/00


Subject: shift summary, owl 02/05/00
From: Valeri Jejer (jejer@uvahea.phys.Virginia.EDU)
Date: Sat Feb 05 2000 - 07:18:23 CST


               Shift Summary for Owl shift, 02/05/00
 
 On Shift: Valeri J., Mark T.
 
 Hours of beam delivered at all intensities: 6.5
 Hours of beam delivered at nominal intensity: 6.5
 Hours of KAMI data: 6.5

 Accounting of lost time:
   Stopping/starting runs, access to roll in TRDs.
 =======================================================================
 
 Runs ended this shift:
 15771 - KAMI_MINBIAS - 11.6k C11 3pi0
 15772 - KTEV_PED
 15773 - KAMI_MUON - >100k evts
 15774 - KAMI_MUON - 260k evts
 15775 - KAMI_MUON - in progress

 ========================================================================
 
 Tapes brought to FCC this shift: none
 
 ========================================================================
 Accesses during shift:
   to move TRDs in.
 
 ========================================================================
 Description of Shift:
run spill time comment
------------------------------------------------------------------------
15771 00:08 RickK. starts KAMI_MINBIAS run with 20" Be.

              00:15 Noticed that NM2S3 has tripped off (SDAQ log shows
                        that this happened at 00:05, so the first 17
                        spills of the run are with the magnet off),
                        call MCR, they turn it back on by 00:32.

              03:17 Stop run after collecting 11.6k 3pi0 in C11. Call
                        MCR and request access to move in TRDs.

              03:35 After access request muon mode with currents for
                        NM2S1=275A and for NM2S2 and NM2S3=0.

15772 03:45 Do short pedestal run to check TRD ADCs.
                        Everything looks OK, so pedestals are not updated.

15773 03:55 Start muon run, stop after 3 spill since B01
                        wasn't sent to file which affected alignment jobs.
                        Modify destination file for KAMI_MUON.

15774 04:10 Start another muon run.

              04:18 Stop the run, update DCGM and TRDG constants.

              04:40 After setting NM4AN to 400A start loooong muon
                        run.

              05:00 Noticed that UCAL16 ADC=0. Since it was bad in
                        a few previous runs, decided not to page anyone.

              06:27 BobT calls and asks to turn off NM3S and NM2SR.
                        MCR tuns off the first, we do the second and set
                        USER LATCH = 1 for these settings. Trigger rate
                        doubled (we had ~3M evts to tape with the magnets
                        on in ~3hrs.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

 * Continue loooooong muon run for two more hours (untill ~9am, as per
   BobT's request), then switch to hadron beam and KAMI_KE3 run; NOTE THAT
   SWEEPERS ARE OFF.

 * update TRDG constants after this muon run (previous update was with
   old DC alignment);



This archive was generated by hypermail 2b27 : Sat Feb 05 2000 - 07:15:39 CST